OpenStudy (anonymous):

let f(x)=x+1 and g(x)=x^2-x. (f+g)(-4)

OpenStudy (anonymous):

are you sure that it's f + g not f OF g?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

well okay, let me try this... So what you do is they are first asking what is G of -4 right? (I'm guessing) so plug in -4 for g(x) so x =-4 therefore it is... -4^2-(-4)=12 so thats your new g... now f do the same f(x)=x+1 4+1 = 5 so F + G 5+12 = 17